# Improving Ground State Accuracy of Variational Quantum Eigensolvers with Soft-coded Orthogonal Subspace Representations

### v1

@@ III.2.3 Relative Gain Factor
In this Section, we introduce and discuss a quantitative metric to assess the improvement in accuracy achieved by the subspace approaches with respect to standard VQE. At first look, one might be tempted to directly use the difference $\delta:=\max_{r}\mathcal{F}_{r,{\rm Algo}}-\max_{r^{\prime}}\mathcal{F}_{r^{\prime},{\rm VQE}}$ between the best fidelities (or the median ones, if we replace $\max$ with $\rm med$}) achieved among different runs $r,r^{\prime}$ for the same system at fixed number of parameters. However, this quantity would value the incremental improvement irrespectively of the absolute value of the fidelities achieved. For example, a $\delta=10\%$ gain obtained when improving the fidelity from $80\%$ to $90\%$ would be considered better than an improvement of $\delta=8\%$ when the fidelity raises from $90\%$ to $98\%$, while intuition seems to suggest to be the opposite case. For this reason, we introduce an empirical measure of the relative gain of a subspace approach with respect to the standard VQE approach, which we call relative gain factor and for which we propose two versions:
@@ III.3 Results for Edwards–Anderson Model
As a more challenging testbed for the subspace approaches investigated, in this Section we consider the 2D Edwards–Anderson spin glass model, described in the introduction to Section III. This model follows the general Ising-like Hamiltonian in Eq. (16). with disordered couplings $J_{ij}$ sampled from independent gaussian distributions with zero mean and unit variance and a uniform transverse field fixed to $h=2$. Every sampling of all the couplings $J_{ij}$ attached to the edges of the $4\times 4$ lattice defines a different realization for the spin glass model. In our numerical investigation, we emulated 10 independent realizations, for each of which we performed 8 independent optimization runs with random initializations of the ansatz parameters in the range $0.2\pi,0.2\pi]$ (same choice as for the TFI model). Since we understood the general behavior of the different approaches from the TFI model in Section III.2, and for the sake of computational accessibility, in this case we focus on a single subspace ansatz configuration, with subspace dimension $K=2$ and $N_{l}=4$ layers for both hard-ortho and soft-ortho. Such choice already provides significant improvements over standard VQE in the TFI case and is found to be reasonably effective also for the EA model (as confirmed below, a posteriori). For the soft-ortho case, we also set $\beta=2.5$, which fulfills weak orthogonality of the lowest energy sector in all the realizations considered.
@@
Let us consider first the evolution of the fidelity $\mathcal{F}_{\rm trc}$ (see definition in Eq. (13) for a reminder) during the optimization for one selected realization of the disordered couplings $J_{ij}$ in Figure 9. Here we observe a non-monotonic behavior of the fidelity with the number of iterations for both standard VQE and hard-ortho approaches, which is even more striking than the one already discussed for the TFI model in Figure 4. The soft-ortho representation, while showing a larger spread across different runs with respect to the TFI case, exhibits instead a more consistent behavior of the fidelities, even if a higher number of iterations (above 1500) could improve even further the final fidelity.
@@
Max (%) Med (%) Max (%) Med (%)
1 0.995 0.417 0.397 7.128
2 3.290 0.934 0.980 0.237
3 0.220 1.614 1.459 2.474
4 0.514 0.623 1.551 5.013
5 0.143 0.594 1.520 0.871
6 0.168 0.528 0.120 0.095
7 0.010 0.397 0.080 0.003
8 0.059 0.070 0.033 2.383
9 0.469 0.333 1.574 5.797
10 0.138 0.617 0.013 0.447
Table 4: Discrepancy $\Delta{\cal F}={\cal F}_{\rm sub}-{\cal F}_{\rm trc}$ between the subspace (Eq. (12)) and $H_{\rm trc}$ fidelity (Eq. (13)) for either best (Max) or median (Med) fidelities across 8 independent runs per 10 realizations of the Edwards–Anderson spin glass model on a $4\times 4$ lattice. The number of layers is fixed to $N_{l}=4$, while the subspace dimension is $K=2$.
As well as for the TFI model, we also monitored the discrepancies between the subspace fidelity $\mathcal{F}_{\rm sub}$ and the $H_{\rm trc}$ fidelity $\mathcal{F}_{\rm trc}$, reporting relevant aggregated data in Table 4. For some realizations, we find somewhat larger discrepancies, possibly due to the increased complexity of the EA model with respect to the TFI one.
@@
Even in these cases, however, looking at final ${\cal F}_{\rm trc}$ fidelities in Table 5, we still observe a significant improvement of the soft-ortho approach with respect to both standard VQE and hard-ortho ones.
@@
To better quantify the improvement over VQE across different disorder realizations, we consider again the relative gain factor introduced in Section III.2.3 and defined in Eqs. (20) and (21). The results are summarized in Table 6, while a more general overview is provided in Figure 10, showing the distribution of relative gain factors for the 10 disorder realizations considered. As observed for the TFI model in Table 3, comparing best runs ($\mathcal{G}_{\min}$) between the hard-ortho representation and standard VQE shows the former not providing any sensible advantage, even underperforming the latter on several realizations. Soft-ortho, instead, shows a clear improvement for all disorder realizations, with a gain factors with medians over the realizations concentrating around 2.5 for both versions ($\mathcal{G}_{\rm med}$ and $\mathcal{G}_{\min}$). For reference, the realization shown in Figure 9 (labeled as ‘10’ in Tables 6) is on the mid-lower side of the gains distribution of Figure 10, with $\mathcal{G}_{\rm med}=1.68$ and $\mathcal{G}_{\min}=2.30$. The relative gain factors achieved by the soft-ortho representation on the EA model on a square $4\times 4$ lattice are lower than those observed for the TFI model in Table 3 on a $3\times 3$ lattice, likely due to the increased complexity of the energy landscape for the former model, which makes the optimization more challenging overall. However, the results shown here confirm the effectiveness of the soft-ortho subspace representation also for slightly larger and more challenging disordered models, where both standard VQE and hard-ortho approaches struggle to provide accurate ground state estimates.
@@ IV Discussion and Conclusions
For both models numerically investigated in this work, we identified a non-monotonic behavior of fidelity metrics during the optimization (see Figures 4 and 9), signaling a possible underfitting phenomenon where the variational ansatz is not expressive enough to represent accurately the ground state but the cost function minimization steers the parameters towards regions of the parameter space where the overlap with eigenstates just above the ground state is favored (see Figure 6). This issue affects both standard VQE and hard-ortho subspace representations, while the soft-ortho approach seems to mitigate it, providing a more consistent convergence behavior and final accuracies. We connect these improvements to the increased expressibility granted by the soft-ortho representation, as argued in Section II.4. For the same reason, the typical circuit depths needed to reach final accuracies comparable or even higher than standard VQE is greatly reduced when using soft-ortho representations, as evident from Figures 7 and 8. In the NISQ era, this is a relevant advantage, since it allows to mitigate the detrimental effects of decoherence by employing relatively shallow circuits to represent the subspace frame states. Even just extending the search space from subspace dimension $K=1$ (standard VQE) to $K=2$ significantly improves the accuracy of the ground state estimate at the end of the optimization and after diagonalization into the optimized subspace. However, we also show that the hard-ortho representation, employed in the SSVQE PhysRevResearch.1.033062 and MCVQE PhysRevLett.122.230401 approaches, does not provide significant improvements over standard VQE in reaching sufficiently high fidelities for the ground state estimates, while the soft-ortho representation systematically outperforms them. On the other hand, increasing the subspace dimension beyond $K=2$ does not seem to improve the fidelity and should be traded off with the total optimization time and the risk of convergence issues within a fixed number of iterations.
@@
As future outlooks, it would be interesting to analyze quantitatively the expressibility of subspace representations, possibly starting from well-known metrics for single-state ansätze from literature ExpressibilitySim2019; Funcke2021dimensional; Arrasmith_2022; PRXQuantum.3.010313; Brozzi:2025omu and extending them to the subspace case. Another possibly relevant aspect to investigate would the effect of sampling noise and quantum noise, affecting independently both the evaluation of the cost function during the optimization and the final diagonalization stage, and impacting on the accuracies of final ground state estimates and the relative gains (defined in Section III.2.3) with respect to standard VQE. Finally, it would be interesting to consider variations of the algorithm employing the soft-ortho representation as described in Section II.1.2, possibly exploring adaptive schemes with dynamic adjustments of the subspace dimension $K$ (i.e., adding/removing states) or varying the orthogonality penalty parameter $\beta$ during the optimization. It is also worth exploring the combination of well-established ansatz adaptive design techniques (such as ADAPT-VQE grimsley2019adaptive) with the soft-coded orthogonal subspace representation introduced here.

### v2

@@ III.2.3 Relative Gain Factor
In this Section, we introduce and discuss a quantitative metric to assess the improvement in accuracy achieved by the subspace approaches with respect to standard VQE. At first look, one might be tempted to directly use the difference $\delta:=\max_{r}\mathcal{F}_{r,{\rm Algo}}-\max_{r^{\prime}}\mathcal{F}_{r^{\prime},{\rm VQE}}$ between the best fidelities (or the median ones, if we replace $\max$ with $\rm med$}) achieved among different runs $r,r^{\prime}$ for the same system at fixed number of parameters. However, this quantity would value the incremental improvement irrespectively of the absolute value of the fidelities achieved. For example, a $\delta=10\%$ gain obtained when improving the fidelity from $80\%$ to $90\%$ would be considered better than an improvement of $\delta=8\%$ when the fidelity raises from $90\%$ to $98\%$, while intuition suggests the opposite. For this reason, we introduce an empirical measure of the relative gain of a subspace approach with respect to the standard VQE approach, which we call relative gain factor and for which we propose two versions:
@@ III.3 Results for Edwards–Anderson Model
As a more challenging testbed for the subspace approaches investigated, in this Section we consider the 2D Edwards–Anderson spin glass model, described in the introduction to Section III. This model follows the general Ising-like Hamiltonian in Eq. (16). with disordered couplings $J_{ij}$ sampled from independent gaussian distributions with zero mean and unit variance and a uniform transverse field fixed to $h=2$. Every sampling of all the couplings $J_{ij}$ attached to the edges of the $4\times 4$ lattice defines a different realization for the spin glass model. In our numerical investigation, we emulated 10 independent realizations, performing 8 independent optimization runs each, with random initializations of the ansatz parameters in the range $0.2\pi,0.2\pi]$ (same choice as for the TFI model). Building on the insights gained from the TFI model in Section III.2, and for the sake of computational accessibility, in this case we focus on a single subspace ansatz configuration, with subspace dimension $K=2$ and $N_{l}=4$ layers for both hard-ortho and soft-ortho. This choice already provides significant improvements over standard VQE in the TFI case and is found to be reasonably effective also for the EA model (as confirmed below, a posteriori). For the soft-ortho case, we set $\beta=2.5$, which fulfills weak pairwise orthogonality of the frame at convergence in all the realizations considered.
@@
Let us consider first the evolution of the fidelity $\mathcal{F}_{\rm trc}$ (see definition in Eq. (13)) during the optimization for one selected realization of the disordered couplings $J_{ij}$ in Figure 9. Here we observe a non-monotonic behavior of the fidelity with the number of iterations for both standard VQE and hard-ortho approaches, which is even pronounced than the one already discussed for the TFI model in Figure 4. The soft-ortho representation, while showing a larger spread across different runs with respect to the TFI case, exhibits instead a more consistent behavior of the fidelities, even if a higher number of iterations (above 1500) could improve even further the final fidelity.
@@
$\max{\Delta{\cal F}}$ % ${\rm med}{\Delta{\cal F}}$ % $\max{\Delta{\cal F}}$ % ${\rm med}{\Delta{\cal F}}$ %
1 0.446 0.25(5) 8.50 3.9(17)
2 1.923 0.55(37) 1.14 0.31(22)
3 1.268 0.90(10) 3.71 0.98(38)
4 0.417 0.31(6) 8.00 2.6(14)
5 1.026 0.57(27) 5.43 1.76(75)
6 0.484 0.30(5) 7.16 <1.5
7 0.295 <0.08 0.52 0.11(7)
8 0.054 0.025(7) 2.20 0.80(48)
9 0.336 0.20(6) 5.41 <2.2
10 0.446 0.20(9) 0.49 0.19(9)
[…]
As with the TFI model, we also monitored the discrepancies between the subspace fidelity $\mathcal{F}_{\rm sub}$ and the $H_{\rm trc}$ fidelity $\mathcal{F}_{\rm trc}$, reporting relevant aggregated data in Table 4. For some realizations, we find somewhat larger discrepancies, possibly due to the increased complexity of the EA model with respect to the TFI one.
@@
However, even in these cases, looking at the final ${\cal F}_{\rm trc}$ fidelities in Table 5, we still observe a significant improvement of the soft-ortho approach with respect to both the standard VQE and the hard-ortho one.
@@
To better quantify the improvement over VQE across different disorder realizations, we consider again the relative gain factor introduced in Section III.2.3 and defined in Eqs. (20) and (21). The results are summarized in Table 6, while a more general overview is provided in Figure 10, showing the distribution of relative gain factors for the 10 disorder realizations considered. As observed for the TFI model in Table 3, regarding best runs ($\mathcal{G}_{\min}$) hard-ortho representation provides no tangible advantage with respect to standard VQE, even underperforming on several disorder realizations. Soft-ortho, instead, shows a clear improvement for all disorder realizations, with median gain factors fluctuating around 2.5 for both versions ($\mathcal{G}_{\rm med}$ and $\mathcal{G}_{\min}$). For reference, the realization shown in Figure 9 (labeled as ‘10’ in Tables 6) is on the mid-lower side of the gains distribution of Figure 10, with $\mathcal{G}_{\rm med}=1.7(2)$ and $\mathcal{G}_{\min}=2.30$. The relative gain factors achieved by the soft-ortho representation on the EA model on a square $4\times 4$ lattice are lower than those observed for the TFI model in Table 3 on a $3\times 3$ lattice, likely due to the increased complexity of the energy landscape of the former model, which makes the optimization more challenging overall. However, the results shown here confirm the effectiveness of the soft-ortho subspace representation also for larger and more challenging disordered models, where both standard VQE and hard-ortho approaches struggle to provide accurate ground state estimates.
@@ IV Discussion and Conclusions
For both models numerically investigated in this work, we identified a non-monotonic behavior of fidelity metrics during the optimization (see Figures 4 and 9), signaling a possible underfitting phenomenon where the variational ansatz is not expressive enough to represent accurately the ground state but the cost function minimization steers the parameters towards regions of the parameter space where the overlap with eigenstates just above the ground state is favored (see Figure 6). This issue affects both standard VQE and hard-ortho subspace representations, while the soft-ortho approach seems to mitigate it, providing a more consistent convergence behavior and final accuracies. We connect these improvements to the increased expressibility granted by the soft-ortho representation, as argued in Section II.4. For the same reason, the typical circuit depths needed to reach final accuracies comparable or even higher than standard VQE are greatly reduced when using soft-ortho representations, as evident from Figures 7 and 8. In the NISQ era, this is a relevant advantage, since it allows mitigating the detrimental effects of decoherence by employing relatively shallow circuits to represent the subspace frame states. Merely extending the search space from subspace dimension $K=1$ (standard VQE) to $K=2$ significantly improves the accuracy of the ground state estimate at the end of the optimization and after diagonalization into the optimized subspace. However, we also show that the hard-ortho representation, employed in the SSVQE PhysRevResearch.1.033062 and MCVQE PhysRevLett.122.230401 approaches, does not provide significant improvements over standard VQE in reaching sufficiently high fidelities for the ground state estimates, while the soft-ortho representation systematically outperforms them. On the other hand, increasing the subspace dimension beyond $K=2$ does not seem to improve the fidelity and should be traded off with the total optimization time and the risk of convergence issues within a fixed number of iterations.
@@
As a future direction, it would be interesting to analyze quantitatively the expressibility of subspace representations, possibly starting from well-known metrics for single-state ansätze from literature ExpressibilitySim2019; Funcke2021dimensional; Arrasmith_2022; PRXQuantum.3.010313; Brozzi:2025omu and extending them to the subspace case. Another possibly relevant aspect to investigate would be the effect of sampling noise and quantum noise, affecting independently both the evaluation of the cost function during the optimization and the final diagonalization stage, and impacting the accuracies of final ground state estimates and the relative gains (defined in Section III.2.3) with respect to standard VQE. Finally, it is possible to consider variations of the algorithm employing the soft-ortho representation as described in Section II.1.2, exploring adaptive schemes with dynamic adjustments of the subspace dimension $K$ (i.e., adding/removing states) or varying the orthogonality penalty parameter $\beta$ during the optimization. It is also worth exploring the combination of well-established ansatz adaptive design techniques (such as ADAPT-VQE grimsley2019adaptive) with the soft-coded orthogonal subspace representation introduced in this work.